River Wild is considering purchasing a water park in Oakland, California​, for $1,950,000.The new facility will generate annual net cash inflows of $495,000 for eight years. Engineers estimate that the facility will remain useful for eight years and have no residual value. The company uses​ straight-line depreciation. Its owners want payback in less than five years and an ARR of 10​% or more. Management uses a 14% hurdle rate on investments of this nature.

Requirements

1.Compute the payback​ period, the​ ARR, the​ NPV, and the approximate IRR of this investment.​ (If you use the tables to compute the​ IRR, answer with the closest interest rate shown in the​ tables.)

2.Recommend whether the company should invest in this projec

urchasing a water park in Oakland, California, for $1,950,000. The new facility will generate annual net cash inflows of $495i Reference Future Value of Annuity of $1 Periods 1% 3% 1.000 2.030 3.091 4.184 5.309 6.468 7.662 8.892 10.159 11.464 4% 1.00t.River Wild is considering purchasing a water park in Oakland, California, for $1,950,000. The new facility will generate annuA Reference Periods 2% 4% 1% 1.000 2.010 3.030 4.060 5.101 6.152 7.214 8.286 9.369 10.462 11.567 12.683 13.809 14.947 16.097
Answer #1

Solution 1:

Payback period = Initial investment / Annual cash inflows = $1,950,000 / $495,000 = 3.94 years

Annual depreciation = $1,950,000 / 8 = $243,750

Annual net income = Annual cash inflows - Depreciation = $495,000 - $243,750 = $251,250

Average investment = (Cost + Residual value) / 2 = ($1,950,000 + 0)/2 = $975,000

ARR = Annual net income / Average investment = $251,250 / $975,000 = 25.77%

Computation of NPV
Particulars Period Amount PV factor at 14% Present Value
Cash outflows:
Initial investment 0 $1,950,000.00 1 $1,950,000
Present Value of Cash outflows (A) $1,950,000
Cash Inflows
Annual cash inflows 1-8 $495,000.00 4.639 $2,296,305
Present Value of Cash Inflows (B) $2,296,305
Net Present Value (NPV) (B-A) $346,305

Cumulative PV factor at IRR = Initial investment / Annual cash inflows = $1,950,000 / $495,000 = 3.939

Refer PV factor table, this factor falls at nearest to IRR = 19%

Solution 2:

Company should invest in this project as all criteria are met.